Circular No. 6786

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

                URL http://cfa-www.harvard.edu/iau/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 1997ef IN UGC 4107 P. Garnavich, S. Jha, R. Kirshner, and P. Challis, Harvard-Smithson-

ian Center for Astrophysics; and D. Balam, University of Victoria, report: “CCD images of this object were obtained with the F. L. Whipple Observatory (FLWO) 1.2-m telescope by W. Brown and C. Briceno, and with the Dominion Astrophysical Observatory 1.8-m Plaskett telescope, yielding the following V magnitude estimates (relative to the star of V = 17.05 that is 1′ to the southwest of UGC 4107): Nov. 28.4 UT, 16.75; 29.5, 16.67; 30.5, 16.57; Dec. 1.3, 16.56; 3.6, 16.55; 12.3, 16.50. The photometry shows a rapid rise in late November, but the brightness has currently leveled out near V = 16.5. Assuming the object is at the distance of UGC 4107 and is unreddened, the absolute V magnitude of the plateau is −17.1 for H_{0} = 65 km s^{−1} Mpc^{−1}. A series of spectra taken with the FLWO 1.5-m telescope between Nov. 26 and Dec. 8 confirms the trend of increasing wavelength for the broad features that was noted on IAUC 6783. The largest peak near 520 nm has shifted by an average of 1500 km/s per day, covering 513 to 540 nm in 10 days. The flux minimum near 600 nm has become deeper and narrower and now resembles the Si II absorption often found in type-I supernovae. The maximum near 440 nm has strengthened and narrowed and currently peaks at an observed wavelength of 450.1 nm. A spectrum (range 320–850 nm), obtained with the Multiple-Mirror Telescope on Dec. 5, shows no strong interstellar absorption lines of Na I or Ca II at the redshift of UGC 4107, to an equivalent-width limit of 0.02 nm, implying little extinction to this object. The spectrum also shows a sharp drop in flux blueward of 375 nm; this may be due to Ca II absorption out to high velocities, or even to a Balmer jump (but no other signature of hydrogen is clearly present).”

COMET 78P/GEHRELS 2 Visual m_{1} estimates: Nov. 4.16 UT, 12.8 (R. J. Bouma, Groningen,

The Netherlands, 0.25-m refl.); 6.05, 12.3 (V. S. Nevski, Vitebsk, Belarus, 0.30-m refl.); Dec. 9.19, 11.6 (A. Pereira, Cabo da Roca, Portugal, 0.25-m refl.); 11.81, 12.2 (A. Pearce, Nedlands, W. Australia, 0.20-m refl.).

GRB 971208 Corrigendum. On IAUC 6785, line 7, for Dec. 8.335 UT read Dec.

8.325 UT

1997 December 12  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1997 CBAT  Daniel W. E. Green